OpenStudy (anonymous):

lim x→0+ sin x/sqrt(x) Hey guys, how do?

OpenStudy (zarkon):

\[\frac{\sin(x)}{\sqrt{x}}=\sqrt{x}\frac{\sin(x)}{x}\]

OpenStudy (anonymous):

Can you elaborate a little? I don't understand why you can just pull the sqrt(x) out and replace it with an x.

OpenStudy (zarkon):

\[x=\sqrt{x}\cdot \sqrt{x}\]

OpenStudy (anonymous):

Ah. Rock n roll man. Thanks.